Exercise 21-7

Rensing Ltd. estimates sales for the second quarter of 2017 will be as follows.

Month

Units

April 2,560
May 2,420
June 2,340


The target ending inventory of finished products is as follows.

March 31 2,040
April 30 2,200
May 31 2,200
June 30 2,300


2 units of material are required for each unit of finished product. Production for July is estimated at 2,800 units to start building inventory for the fall sales period. Rensing’s policy is to have an inventory of raw materials at the end of each month equal to 50% of the following month’s production requirements.

Raw materials are expected to cost $7 per unit throughout the period.

Calculate the May raw materials purchases in dollars.

Raw material purchases cost $
Answer #1

Production in May = 2420+2200-2200 = 2420

Production in June = 2340+2300-2200 = 2440

Raw material purchase = (2420*2)+(2440*50%)-(2420*2*50%) = 4860

Raw material purchase = 4860*7 = 34020